OpenStudy (anonymous):

Which set of line segments with the given lengths can be formed into a right triangle? A) 5, 6, 11 B) 5, 9, 10 C) 5, 13, 18 D) 5, 12, 13

hero (hero):

Hint: 5, 12, 13 is a known pythagorean triple since 5^2 + 12^2 = 13^2 25 + 144 = 169

OpenStudy (anonymous):

So it wpuld have to be a triple pythagorean?

OpenStudy (anonymous):

@Hero

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

so its D?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw